# Enabling Robust Exoplanet Atmospheric Retrievals with Gaussian Processes

### v1

@@ 5.3 Concluding Remarks
Y.R. thanks Dr. Aishwarya Iyer for meaningful conversations during the early stages of this project. This research was enabled thanks to support from NASA XRP grant 80NSSC24K0160. L.W. thanks the Heising-Simons Foundation for their funding through the 51 Pegasi b Postdoctoral Fellowship, and Arizona State University for their support through the Presidential Postdoctoral Fellowship Program. Y.R., L.W., and M.R.L. acknowledge Research Computing at Arizona State University for providing HPC and storage resources that have significantly contributed to the research results reported within this manuscript. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. The document release number is LLNL-JRNL-2003715.
@@ Appendix A Synthetic Model with a Higher-Amplitude Feature
In this case, the global GP is again preferred over a traditional retrieval, this time by 17.3; however, the combined local and global retrieval is preferred over the global retrieval by 7.1, implying a strong need for a local kernel. Thus, we conclude that the lack of strong “detection” of the local kernel in Section 3 is driven by the small amplitude of the feature, rather than the inefficiency of a local kernel. The retrieved spectra, absorber abundances, and hyperparameters are shown in Figure A1.
Appendix B Synthetic Model GP Contribution Functions
In Figure A2, we show the contribution functions of the GP for the synthetic spectrum of WASP-96 b (Section 3). This visually isolates the wavelengths at which the GP’s local kernels are acting. Negative values correspond to regions where the GP had to shift the mean spectrum down to match data, while positive values correspond to a shift up. We note that for the traditional framework, the GP contribution function is a constant value of zero, as expected in the case without a GP. The local kernel only case is able to identify the injected feature at 1.65 m, while the global-only case inflates the uncertainty of the entire spectrum. By combining both, the GP is able to account for regions of high correlation without overcorrecting across the entire spectrum.
Appendix C Synthetic Model GP Contribution Functions
@@
Appendix D WASP-96 b GP Contribution Functions
Here, we show the contribution functions of the GP for the NIRISS/SOSS spectrum of WASP-96 b in Figure A3. Again, in the traditional case, the GP contribution function is zero. The two regions identified by the GP are at 1.18 m (below the spectrum) and m (above the spectrum). The latter region affects the CO2 abundance, which is constrained largely by fitting the absorption feature at m.
D.1 WASP-96 b - Retrieved Cloud and Aerosol Properties
The potential for clouds and hazes of WASP-96 b’s atmosphere has been extensively studied (Nikolov et al., 2018; Yip et al., 2020; Nikolov et al., 2022b; McGruder et al., 2022; Radica et al., 2023; Taylor et al., 2023; Samra et al., 2023; Zamyatina et al., 2024). While GCMs show the possibility for a cloudy terminator, observations from VLT/FORS2 and Magellan/IMACS find a largely cloud-free terminator, inferred from the width of the Na doublet feature. The cloud-free nature of WASP-96 b has also been inferred from HST and JWST observations. However, analysis of the JWST observations (Taylor et al., 2023) implied a strong scattering slope from aerosols in the atmosphere, which was not identified in optical data from HST and ground-based instruments. This slope was found to be degenerate with the Na abundance, as the Na doublet feature at m could act as a scattering slope in NIRISS/SOSS spectra, as only the red wing of the feature is observed.
We are able to reproduce and confirm previous results from Taylor et al. (2023), which found no need for a grey cloud deck (, bar) but a strong need for a haze-enhanced scattering slope (, ). We also find a degeneracy between haze parameters and cloud fraction, implying that stronger scattering slopes are more consistent with a lower cloud fraction. Notably, the GP-aided retrievals retrieve a lower cloud fraction than those found by the traditional retrievals (Table 2), as well as a stronger scattering slope amplitude that is negatively degenerate with cloud fraction; this, combined with the relative lack of a scattering slope in the NIRISS data, implies that only a few data points towards the blue edge of the spectrum are likely driving the need for a strongly haze-enhanced slope; this is also noted in Taylor et al. (2023), who found a degeneracy between the Na wing at the blue edge of the bandpass and the haze scattering slope. The retrieved cloud and haze properties from our models are shown in Figure A4.
Appendix E Retrievals of WASP-96 b without Background Contamination Correction
Here, we compare our results from retrievals of the case of data with background contamination to the case of data with contamination corrected out. The retrieved spectra and abundances can be seen in figure A5. We find that there is a stronger need for a GP in the case of background contamination (G, G1L, and G2L cases preferred over a traditional retrieval by 6.1, 6.3, and 6.2, respectively). However, the local kernels do not identify the background contamination regions (1.3–1.7 and 2.1 m), implying that their effect on the spectrum is minimal (Figure A6).
This leads to two key differences between the datasets; firstly, the amplitude of the global correlation is higher in the uncorrected dataset that contains the contaminants (102 ppm versus 86 ppm in the corrected case) in order to marginalize over the 2.1 m contamination feature with the global kernel. Secondly, due to this higher amplitude in the uncorrected case, the feature at 1.18 m is also marginalized over with the global kernel, and does not require a local kernel to be placed, whereas a local kernel is needed to characterize it in retrievals of the corrected data. As such, both local kernels identify the region at m instead.

### v2

@@ 5.3 Concluding Remarks
We thank the referee for thoughtful comments that significantly improved this work. Y.R. thanks Dr. Aishwarya Iyer for meaningful conversations during the early stages of this work. This research was enabled thanks to support from NASA XRP grant 80NSSC24K0160. L.W. thanks the Heising-Simons Foundation for their funding through the 51 Pegasi b Postdoctoral Fellowship, and Arizona State University for their support through the Presidential Postdoctoral Fellowship Program. Y.R., L.W., and M.R.L. acknowledge Research Computing at Arizona State University for providing HPC and storage resources that have significantly contributed to the research results reported within this manuscript. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344. The document release number is LLNL-JRNL-2003715.
@@ Appendix A Synthetic Model with a Higher-Amplitude Feature
In this case, the global GP is again preferred over a traditional retrieval, this time by 17.3 (); however, the combined local and global retrieval is preferred over the global retrieval by 7.1 (), implying a strong need for a local kernel. Thus, we conclude that the lack of strong “detection” of the local kernel in Section 3 is driven by the small amplitude of the feature, rather than the inefficiency of a local kernel. The retrieved spectra, absorber abundances, and hyperparameters are shown in Figure A1.
Appendix B WASP-96 b Retrieved Parameters